# Nonlinearly Preconditioned Gradient Methods under Generalized Smoothness

### v1

@@ 2 The extended anisotropic descent inequality
In this section we extend the definition of anisotropic smoothness from [laude2022anisotropic] to our setting where $\phi$ is potentially nonsmooth and provide sufficient conditions for a smooth function $f$ to satisfy this generalized descent inequality. The proofs can be found in Appendix B.
@@
$f(x)$ $\leq f(\bar{x})+\bar{L}\left[(L^{-1}\star\phi)(x-\bar{y})-(L^{-1}\star\phi)(\bar{x}-\bar{y})\right],$ (4)
@@
Note that inequality (4) is well-defined for $\phi$ without full domain, but does not provide information for points $x\in\mathbb{R}^{n}$ such that $L(x-\bar{y})\notin\operatorname{dom}\phi$. The intuition behind this extension of [laude2022anisotropic] is straightforward: we allow for two different smoothness constants that play a complementary role. It can be checked that for $\operatorname{dom}\phi=\mathbb{R}^{n}$, Definition 2.1 reduces to [laude2022anisotropic, Definition 3.1] but w.r.t. $\bar{L}\phi$.
Our first result is an extension of the envelope representation of $f$ under anisotropic smoothness [laude2022anisotropic, Proposition 4.1] to our setting where $\phi$ possibly does not have full domain.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ satisfy Definition 2.1. Then, $f(x)=\inf_{y\in\mathbb{R}^{n}}\bar{L}(L^{-1}\star\phi)(x-y)+\xi(y)$ for some proper $\xi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. If, moreover, $f$ is convex, then $f^{*}=\psi+L^{-1}(\bar{L}\star\phi^{*})$ for some lsc and convex $\psi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. This implies that $f^{*}-L^{-1}(\bar{L}\star\phi^{*})$ is convex.
Note that in the convex case Proposition 2.2 describes a conjugate duality between anisotropic smoothness and Bregman strong convexity, parallel to the classical duality between $L$-smoothness and $L^{-1}$-strong convexity. This result along with the envelope representation of $f$ will be utilized later on in order to describe the convergence of the method in the convex setting.
In this paper we are mostly interested in functions that are not covered by the classical $L$-smoothness assumption and thus study reference functions $\phi$ that generate a less restrictive descent inequality. Therefore, we next show that for strongly convex reference functions, the class of anisotropically smooth functions is at least as large as that of Lipschitz smooth ones.
@@ 2.1 Second-order sufficient conditions
In contrast to Euclidean or Bregman smoothness, anisotropic smoothness cannot in general be directly obtained via a second-order condition. This fact becomes apparent by noting that (4) is equivalent to $\bar{x}\in\operatorname*{arg\,min}_{x}g(x)\coloneq\bar{L}(L^{-1}\star\phi)(x-\bar{y})-f(x)$: the second-order necessary condition for the minimality of $\bar{x}$ under Assumption 1.2 is $\bar{L}L\nabla^{2}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x})))-\nabla^{2}f(\bar{x})\succeq 0$, which does not normally imply the convexity of $g$. From the implicit function theorem, the above expression can be written as $\bar{L}L[\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x}))]^{-1}-\nabla^{2}f(\bar{x})\succeq 0$, which is the form that we consider throughout the paper. This condition becomes sufficient when $f,\phi\in\mathcal{C}^{2}(\mathbb{R}^{n})$ are Legendre through [laude2022anisotropic, Proposition 4.1]. Harnessing the connection with $\Phi$-convexity, it is sufficient for general $f$ when $\phi$ is a regular optimal transport cost in light of [Vil08, Theorem 12.46]. Nevertheless, the regularity of $\phi$ is in general hard to verify, since its equivalent form requires the computation of fourth-order derivatives [Vil08, Definition 12.27], and quite restrictive, not holding for many interesting functions. We thus follow a different strategy and study the minimization of $g$ using tools from optimization and nonsmooth analysis.
@@
$\lambda_{\rm max}(H\nabla^{2}f(x))\leq L\bar{L}.$ (5)
In particular, under Assumption 1.2, this reduces to $\lambda_{\rm max}(\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x))\leq L\bar{L}$.
The generalized Jacobian considered in the definition above can in fact be computed for many interesting reference functions. As an example we study the inequality (5) for the reference function of Example 1.7 in subsection D.3. Nevertheless, under Assumption 1.2 we obtain a condition that is easier to check.
@@
Let $\phi$ satisfy Assumption 1.2 and $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$. Then, $f$ satisfies Definition 2.4 if and only if for all $x\in\mathbb{R}^{n}$
$\nabla^{2}f(x)\preceq L\bar{L}[\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1}.$ (6)
@@
$\lambda_{\max}(\nabla^{2}f(x))\leq L\bar{L}\lambda_{\min}([\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1}).$ (7)
Lemma 2.5 as well as the following proposition which describes the monotonicity of the forward operator $T_{{\gamma},{\bar{L}^{-1}}}$ under Definition 2.4 are inspired by [villani2021topics, Exercise 2.57], underlying once again the strong connection of anisotropic smoothness with $\Phi$-convexity. In Proposition 2.6, we show that for $f$ satisfying Definition 2.4, we can make $T_{{\gamma},{\bar{L}^{-1}}}$ a strongly monotone operator by choosing suitable $\gamma>0$.
Proposition 2.6 (nonsmooth forward step monotonicity).
Let $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ satisfy Definition 2.4. Then, the following inequality holds:
$\langle T_{\gamma,\bar{L}^{-1}}(x)-T_{\gamma,\bar{L}^{-1}}(\bar{x}),x-\bar{x}\rangle\geq(1-\gamma L)\|x-\bar{x}\|^{2}$ (8)
for all $x,\bar{x}\in\mathbb{R}^{n}$ and $\gamma>0$. In particular, the forward operator $T_{\gamma,\bar{L}^{-1}}$ is strongly monotone with parameter $1-\gamma L$ for $\gamma<L^{-1}$. Furthermore, under Assumption 1.2 the reverse implication holds: monotonicity of $T_{L^{-1},\bar{L}^{-1}}$ implies $\lambda_{\rm max}(\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x))\leq L\bar{L}$.
Proposition 2.6 allows us to obtain novel insights into the class of $\mathcal{C}^{2}(\mathbb{R}^{n})$ $(L_{0},L_{1})$-smooth functions.
@@
Let $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ be $(L_{0},L_{1})$-smooth. Then,
$T_{{L_{1}^{-1}},{L_{1}/L_{0}}}(x)=x-\tfrac{\delta}{L_{0}+L_{1}\|\nabla f(x)\|}\nabla f(x)$
is monotone for $\delta=1$ and strongly monotone for $10<\delta<$ 1.
@@
Considering the forward operator as defined in Corollary 2.7, the main iteration of the algorithm (2) we analyze becomes:
@@
Note that in this case, [gorbunov2024methods, Algorithm 1] can be viewed as (2) with a conservative choice of $\delta\leq 0.57$. In a similar manner, the stepsize introduced in [vankov2024optimizing, Equation (3.5)] can be viewed as a less tight version of Corollary 2.7. This algorithm is in fact generated by $\phi=h\circ\|\cdot\|$ with $h(x)=-|x|-\ln(1-|x|)$.
Although Corollary 2.7 establishes a connection between $(L_{0},L_{1})$-smoothness and $(L,\bar{L})$-anisotropic smoothness, the simplification used in the proof of the second-order condition is quite restrictive. We provide examples where we can obtain tighter constants utilizing directly Definition 2.4 in Appendix D.
Having obtained a sufficient condition for the (strong) monotonicity of the forward operator in Proposition 2.6, we now move on to providing sufficient conditions for $f$ to be $(L,\bar{L})$-anisotropically smooth.
@@
Let $f\in\mathcal{C}(\mathbb{R}^{n})$ and $T_{{L^{-1}},{\bar{L}^{-1}}}$ be strictly monotone. Let moreover either $\operatorname{dom}\phi$ be bounded or $\operatorname{dom}\phi=\mathbb{R}^{n}$ and the following growth assumption hold:
@@
for all $x\in\mathbb{R}^{n}$ and some $r>0$, $r<L,\beta\in\mathbb{R}$. Then, $f$ is $(L,\bar{L})$-anisotropically smooth relative to $\phi$.
@@ 2.2 How to compute the second-order condition
In this subsection we demonstrate how the second-order condition of Definition 2.4 can be computed for the two different instances of our preconditioned scheme. We consider kernel functions that satisfy Assumption 1.1 and Assumption 1.2 implying that the results of Lemma 1.3 hold.
Separable reference functions. In this case, the condition is simple to compute, since $[\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1}$ is just a diagonal matrix with elements $\alpha_{ii}$ given by
@@
Isotropic reference functions. As already mentioned in the introduction, in the isotropic case $\phi=h\circ\|\cdot\|$ the differentiability of $\phi^{*}$ does not hold for general kernel functions $h$. In the setting considered in this subsection though, we have $\nabla\phi^{*}(y)={h^{*}}^{\prime}(\|y\|)\overline{\operatorname{sgn}}(y)$ with $\phi^{*}\in\mathcal{C}^{2}(\mathbb{R}^{n})$ and the Hessian is given by
$[\nabla^{2}\phi^{*}(y)]^{-1}=\frac{1}{{h^{*}}^{\prime\prime}(\|y\|)}\frac{yy^{\top}}{\|y\|^{2}}+\frac{\|y\|}{{h^{*}}^{\prime}(\|y\|)}\left(I-\frac{yy^{\top}}{\|y\|^{2}}\right)$
@@ 3.1 Nonconvex setting
The convergence of the method to stationary points, in the composite case with an additional nonconvex, nonsmooth term, was established in [laude2022anisotropic, Theorem 5.3] with a fixed stepsize $\gamma\leq L^{-1}$ and in [laude2022anisotropic, Theorem 5.5] using an adaptive linesearch strategy for choosing $\gamma$, always under the assumption that $\phi$ is of full domain. In the smooth setting studied in this paper, we can improve upon the aforementioned results and show that stepsizes $\gamma$ up to $2L^{-1}$ can actually be used in the algorithm.
@@
Let $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ be the sequence of iterates generated by (2) with $\gamma=\alpha L^{-1}$, $\alpha\in(0,2)$, $\lambda=\bar{L}^{-1}$ and let $\beta=|1-\alpha|$. Then we have the following rate:
$\min_{0\leq k\leq K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L(f(x^{0})-f^{\star})}{\bar{L}(1-\beta)(K+1)}.$
@@
Let $\phi=\cosh\circ\|\cdot\|-1$ and $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ be the sequence of iterates generated by (2) with $\gamma=\alpha L^{-1}$, $\alpha\in(0,2)$, $\lambda=\bar{L}^{-1}$ and let $\beta=|1-\alpha|$. Then, the following holds:
$\min_{0\leq k\leq K}\|\nabla f(x^{k})\|\leq\sqrt{\frac{2L\bar{L}(f(x^{0})-f^{\star})}{(1-\beta)(K+1)}}+\frac{L(f(x^{0})-f^{\star})}{(1-\beta)(K+1)}.$

### v2

@@ 2 The extended anisotropic descent inequality
In this section we extend the definition of anisotropic smoothness from (Laude & Patrinos, 2025) to our setting where $\phi$ is potentially nonsmooth and provide sufficient conditions for a smooth function $f$ to satisfy this generalized descent inequality. The proofs can be found in Appendix B.
@@
$f(x)$ $\leq f(\bar{x})+\bar{L}\left[(L^{-1}\star\phi)(x-\bar{y})\right.$ (4)
$\left.-(L^{-1}\star\phi)(\bar{x}-\bar{y})\right],$
@@
Note that inequality (4) is well-defined for $\phi$ without full domain, but does not provide information for points $x\in\mathbb{R}^{n}$ such that $L(x-\bar{y})\notin\operatorname{dom}\phi$. The intuition behind this extension of (Laude & Patrinos, 2025) is straightforward: we allow for two different smoothness constants that play a complementary role, while allowing $\operatorname{dom}\phi\neq\mathbb{R}^{n}$ leads to a more general descent inequality. It can be checked that for $\operatorname{dom}\phi=\mathbb{R}^{n}$, Definition 2.1 reduces to (Laude & Patrinos, 2025, Definition 3.1) but w.r.t. $\bar{L}\phi$.
Our first result is an extension of the envelope representation of $f$ under anisotropic smoothness (Laude & Patrinos, 2025, Proposition 4.1) to our setting where $\phi$ possibly does not have full domain.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ satisfy Definition 2.1. Then, $f(x)=\inf_{y\in\mathbb{R}^{n}}\bar{L}(L^{-1}\star\phi)(x-y)+\xi(y)$ for some proper $\xi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. Moreover, $f^{*}=\psi+L^{-1}(\bar{L}\star\phi^{*})$ for some lsc and convex $\psi:\mathbb{R}^{n}\to\overline{\mathbb{R}}$, implying that $f^{*}-L^{-1}(\bar{L}\star\phi^{*})$ is convex.
Note that Proposition 2.2 describes one direction of a conjugate duality between anisotropic smoothness and Bregman (relative) strong convexity (Lu et al., 2018, Definition 1.2). This result along with the envelope representation of $f$ will be utilized later on in order to describe the convergence of the method in the convex setting.
In this paper we are mostly interested in cost functions that are not covered by the classical $L$-smoothness assumption and thus study reference functions $\phi$ that generate a less restrictive descent inequality. Therefore, we next show that for strongly convex reference functions, the class of anisotropically smooth functions is at least as large as that of Lipschitz smooth ones.
@@ 2.1 Second-order sufficient conditions
In contrast to Euclidean or Bregman smoothness, anisotropic smoothness cannot in general be directly obtained via a second-order condition. This fact becomes apparent by noting that (4) is equivalent to $\bar{x}\in\operatorname*{arg\,min}_{x}g(x)\coloneq\bar{L}(L^{-1}\star\phi)(x-\bar{y})-f(x)$. The second-order necessary condition for the minimality of $\bar{x}$ under Assumption 1.2 then becomes $\bar{L}L\nabla^{2}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x})))-\nabla^{2}f(\bar{x})\succeq 0$, which does not normally imply the convexity of $g$. From the implicit function theorem, the above expression can be written as $\bar{L}L[\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(\bar{x}))]^{-1}-\nabla^{2}f(\bar{x})\succeq 0$, which is the form that we consider throughout the paper. This condition becomes sufficient when $f,\phi\in\mathcal{C}^{2}(\mathbb{R}^{n})$ are Legendre through (Laude & Patrinos, 2025, Proposition 4.1). Harnessing the connection with $\Phi$-convexity, it is sufficient for general $f$ when $\phi$ is a regular optimal transport cost in light of (Villani, 2008, Theorem 12.46). Nevertheless, the regularity of $\phi$ is in general hard to verify, since its equivalent form requires the computation of fourth-order derivatives (Villani, 2008, Definition 12.27), and quite restrictive, not holding for many interesting functions. We thus follow a different strategy and study the minimization of $g$ using tools from optimization and nonsmooth analysis.
@@
$\lambda_{\rm max}(H\nabla^{2}f(x))<L\bar{L}$ (5)
and $\lim_{\|x\|\to\infty}\|T_{{L^{-1}},{\bar{L}^{-1}}}(x)\|=\infty.$ In particular, (5) reduces to $\lambda_{\rm max}(\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))\nabla^{2}f(x))<L\bar{L}$ under Assumption 1.2.
Note that, since $\phi^{*}$ is Lipschitz smooth and convex, from (Hiriart-Urruty et al., 1984, Example 2.2) we know that $H$ is always a positive semi-definite matrix and thus from (Horn & Johnson, 2012, Theorem 1.3.22) with $A=H^{1/2}\nabla^{2}f(x)$ and $B=H^{1/2}$ we have that $H\nabla^{2}f(x)$ has real eigenvalues.
The generalized Jacobian considered in the definition above can in fact be computed for many interesting reference functions. As an example we study the inequality (5) for the reference function of Example 1.7 in subsection D.2. The coercivity assumption on the forward operator in Definition 2.4 is very mild. For example consider a reference function $\phi$ with $\operatorname{dom}\phi\subseteq\overline{\mathbb{B}}(0,1)$ as the isotropic ones generated by the kernels in the four last rows of Table 1. Then, by standard convex conjugacy, $\|\nabla\phi^{*}\|\leq 1$ and $\lim_{\|x\|\to\infty}\|T_{{\gamma},{\lambda}}(x)\|=\infty$ always. We provide more results regarding the norm-coercivity property of $T_{{L^{-1}},{\bar{L}^{-1}}}$ in subsection D.1. It is important to note that when the matrix $H\nabla^{2}f(x)$ is symmetric, we can remove this extra condition on the forward operator, as we show in Appendix D.
Under Assumption 1.2 we obtain a condition that is more straightforward to check.
@@
Let $\phi$ satisfy Assumption 1.2 and $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$. Then, (5) holds if and only if for all $x\in\mathbb{R}^{n}$
$\nabla^{2}f(x)\prec L\bar{L}[\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1}.$ (6)
@@
$\lambda_{\max}(\nabla^{2}f(x))<L\bar{L}\lambda_{\min}([\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1}).$ (7)
Next, we show that under Definition 2.4 the forward operator is actually a global homeomorphism (Facchinei & Pang, 2003, Definition 2.1.9), which we further utilize later on in proving our main result regarding the sufficiency of the second-order condition, Proposition 2.9.
Proposition 2.6.
Let $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ satisfy Definition 2.4. Then, $T_{{L^{-1}},{\bar{L}^{-1}}}$ is injective.
As a byproduct of our analysis, we obtain novel insights into the class of $\mathcal{C}^{2}(\mathbb{R}^{n})$ $(L_{0},L_{1})$-smooth functions.
@@
Let $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ be $(L_{0},L_{1})$-smooth. Then, $T_{{\delta L^{-1}},{\bar{L}^{-1}}}=x-\tfrac{\delta}{L_{0}+L_{1}\|\nabla f(x)\|}\nabla f(x)$ is monotone for $\delta=1$ and strongly monotone for $10<\delta<$ 1.
@@
Considering the forward operator as defined in Corollary 2.7, the main iteration (2) becomes:
@@
Note that in this case, (Gorbunov et al., 2024, Algorithm 1) can be viewed as (2) with a conservative choice of $\delta\leq 0.57$. This algorithm is in fact generated by $\phi=h\circ\|\cdot\|$ with $h(x)=-|x|-\ln(1-|x|)$.
Although Corollary 2.7 establishes new characterizations for $(L_{0},L_{1})$-smoothness, the simplification used in the proof of the second-order condition is quite restrictive. We provide examples where we can obtain tighter constants utilizing directly Definition 2.4 in Appendix D.
Having obtained a sufficient condition for the injectivity of the forward operator in Proposition 2.6, we now move on to providing sufficient conditions for $f$ to be $(L,\bar{L})$-anisotropically smooth.
@@
Let $f\in\mathcal{C}^{1}(\mathbb{R}^{n})$ and $T_{{L^{-1}},{\bar{L}^{-1}}}$ be injective. Let moreover either (i) $\operatorname{dom}\phi$ be bounded or (ii) $\operatorname{dom}\phi=\mathbb{R}^{n}$ and the following growth condition hold
@@
for all $x\in\mathbb{R}^{n}$ and some $r,\beta\in\mathbb{R}$ such that $0<r<L$. Then, $f$ is $(L,\bar{L})$-anisotropically smooth relative to $\phi$.
@@
By combining Corollary 2.7 and Proposition 2.9 we can easily obtain the following result that describes the relation between $(L,\bar{L})$-anisotropic smoothness and $(L_{0},L_{1})$-smoothness.
Corollary 2.11.
Let $f\in\mathcal{C}^{2}(\mathbb{R}^{n})$ be $(L_{0},L_{1})$-smooth. Then, $f$ is $(\delta L_{1},L_{0}/L_{1})$-anisotropically smooth relative to $\phi(x)=-\|x\|-\ln(1-\|x\|)$ with $\delta\in(0,1)$.
@@ 2.2 How to compute the second-order condition
In this subsection we demonstrate how the second-order condition of Definition 2.4 can be computed for the two different instances of our preconditioned scheme. We consider kernel functions that satisfy Assumptions 1.1 and 1.2 implying that the results of Lemma 1.3 hold, lifting thus the need to compute generalized Jacobians.
Anisotropic reference functions. Thanks to the separability of $\phi$, the condition is simple to compute, since $[\nabla^{2}\phi^{*}(\bar{L}^{-1}\nabla f(x))]^{-1}$ is just a diagonal matrix with elements $\alpha_{ii}$ given by
@@
Isotropic reference functions. As already mentioned in the introduction, in the isotropic case $\phi=h\circ\|\cdot\|$ the differentiability properties of $\phi^{*}$ depend on those of $h^{*}$. In the setting considered in this subsection though, we have $\nabla\phi^{*}(y)={h^{*}}^{\prime}(\|y\|)\overline{\operatorname{sgn}}(y)$ with $\phi^{*}\in\mathcal{C}^{2}(\mathbb{R}^{n})$ and the Hessian is given by
$[\nabla^{2}\phi^{*}(y)]^{-1}=\frac{1}{{h^{*}}^{\prime\prime}(\|y\|)}\frac{yy^{T}}{\|y\|^{2}}+\frac{\|y\|}{{h^{*}}^{\prime}(\|y\|)}\left(I-\frac{yy^{T}}{\|y\|^{2}}\right)$
@@ 3 Algorithmic analysis
Figure 2: Minimizing $\tfrac{1}{4}\|x\|^{4}$ using (2). The figure on the left corresponds to $\phi_{1}(x)=\cosh(\|x\|)-1$, the middle one to $\phi_{2}(x)=\exp(\|x\|)-\|x\|-1$ and the one on the right to $\phi_{3}(x)=-\|x\|-\ln(1-\|x\|)$. We choose values of $\bar{L}$, set $\lambda=\bar{L}^{-1}$ and then compute $\gamma=L^{-1}$ with $L$ as in Appendix D.
@@ 3.1 Nonconvex setting
The convergence of the method to stationary points, in the composite case with an additional nonconvex, nonsmooth term, was established in (Laude & Patrinos, 2025, Theorem 5.3) with a fixed stepsize $\gamma\leq L^{-1}$ and in (Laude & Patrinos, 2025, Theorem 5.5) using an adaptive linesearch strategy for choosing $\gamma$, always under the assumption that $\phi$ is of full domain. In the smooth setting studied in this paper, where $\phi$ is also even, we can improve upon the aforementioned results and show that stepsizes $\gamma$ up to $2L^{-1}$ can actually be used in the algorithm.
@@
Let Assumption 3.1 hold and $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ be the sequence of iterates generated by (2) with $\gamma=\alpha L^{-1}$, $\alpha\in(0,2)$, $\lambda=\bar{L}^{-1}$ and let $\beta=1-|1-\alpha|$. Then we have the following rate:
$\min_{0\leq k\leq K}\phi(\nabla\phi^{*}(\bar{L}^{-1}\nabla f(x^{k})))\leq\frac{L(f(x^{0})-f^{\star})}{\bar{L}\beta(K+1)}.$
@@
Let Assumption 3.1 hold, $\phi=\cosh\circ\|\cdot\|-1$ and $\{x^{k}\}_{k\in\mathbb{N}_{0}}$ be the sequence of iterates generated by (2) with $\gamma=\alpha L^{-1}$, $\alpha\in(0,2)$, $\lambda=\bar{L}^{-1}$ and $\beta=1-|1-\alpha|$. Then, the following holds for $P_{0}=f(x^{0})-f^{\star}$:
$\min_{0\leq k\leq K}\|\nabla f(x^{k})\|\leq\sqrt{\frac{2L\bar{L}P_{0}}{\beta(K+1)}}+\frac{LP_{0}}{\beta(K+1)}.$